OpenStudy (anonymous):

Write the quadratic equation in general form. What is the value of b√2 - 4ac? 1 = 2x√2 + 5x? 1) 17 2) 33 3) 34

OpenStudy (mathmale):

Howdy, Dylan! This problem requires several steps, so please don't try to solve it in one shot. First of all, what does a quadratic in standard form look like? Secondly: can you write the given equation in standard form?

OpenStudy (anonymous):

Hi (: And I honestly have no idea what any of this even means :(

OpenStudy (mathmale):

Just a thought: Are you positive that you've copied the question correctly? 1 = 2x√2 + 5x can be put into standard form for a quadratic equation, but the inclusion of that square root symbol is highly unusual.

OpenStudy (mathmale):

Dylan: Come to think of it, perhaps this equation, as you've typed it, isn't a quadratic equation at all. A quadratic equation in standard form looks like this: \[ax ^{2}+bx+c=0.\]

OpenStudy (mathmale):

Your 1 = 2x√2 + 5x has only the first power of x, not the second. Dylan: Come on, let's continue. I'm positive you can learn this stuff!

OpenStudy (anonymous):

Write the quadratic equation in general form. What is the value of b^2 - 4ac? 1 = 2x^2 + 5x? ... My bad /_.

OpenStudy (mathmale):

Aha. I was growing more and more positive that something was not quite right. To get you off on the right track: Let's re-write this equation in standard form as explained a few minutes earlier.

OpenStudy (anonymous):

Ok...

OpenStudy (mathmale):

1 = 2x^2 + 5x becomes 2x^2 + 5x - 1 = 0, which corresponds with ax^2 + bx + c = 0. Sound OK to you?

OpenStudy (mathmale):

Please compare these two equations, and then identify the values of a, b and c. For example, a=2. What's b? What's c?

OpenStudy (anonymous):

I wasn't taught this... I was in IMP where i had a terrible teacher... That's why I failed... I'm taking credit recovery now and they gave me algebra... So I don't get this not even the slightest bit .-.

OpenStudy (mathmale):

Sounds like a tough situation, Dylan, but there are a good number of people here on OpenStudy who really do care and will help you if you stick with us. We're talking about quadratic equations here. The general form of a quadratic equation in standard form is ax^2 + bx + c = 0.

OpenStudy (mathmale):

Our quadratic equation is 2x^2 + 5x - 1 = 0.

OpenStudy (mathmale):

Compare the coefficients of the x^2 terms. In the standard form equation, this coeff. is a. In our quadratic, this coeff. is 2. Therefore, a=2. Likewise, the coeff. of x in the standard form equation is b; in our quadratic, it's 5. Therefore, b=5. Please find c.

OpenStudy (anonymous):

C=x^2?

OpenStudy (mathmale):

D: Please compare the last term of the quad. eqn. in std. form to the last term in OUR quad. eqn. Then summarize what you see by writing c = (appropriate value).

OpenStudy (anonymous):

I'm so confused /.\

OpenStudy (mathmale):

We'll get back to this. For now, please take my word for this: a=2, b=5, and c=-1.

OpenStudy (mathmale):

Dylan, this problem asks you to evaluate (find the value of) \[b ^{2}-4ac,\]

OpenStudy (mathmale):

which we call the "discriminant". If a=2, b=5 and c=-1, what is the value of this discriminant? \[b ^{2}=5^{2}=25,4ac=4(2)(-1).\]

OpenStudy (mathmale):

Then, \[-4ac=-4(2)(-1)=8.\] Finally, \[b ^{2}-4ac=25+8=33.\]

OpenStudy (mathmale):

That's it. Dylan: I'm sure you understand some of this, even if you don't feel comfortable with the whole procedure. Please tell me which parts are clear for you and which are not.

OpenStudy (mathmale):

Dylan, would you mind typing out one or two more of the homework problems you're supposed to do? I'd like to get a feeling for how challenging your course is.